OpenStudy (mckenzieandjesus):

Find a solution to the following system of equations. mc002-1.jpg mc002-2.jpg A. (-8, -15) B. (0,1) C.(-2, -15) D.(2,5)

OpenStudy (mckenzieandjesus):

-5x+y=-5 -4x+2y=2 for the images

OpenStudy (yolomcswagginsggg):

@IrishBoy123 @Jemurray3 @wio @Broskishelleh

OpenStudy (yolomcswagginsggg):

this guy needs halp ;p

OpenStudy (yolomcswagginsggg):

@haleyelizabeth2017

OpenStudy (haleyelizabeth2017):

Plug in y=5x-5 (-5x+y=-5 add 5x to both sides) for y in the second equation: -4x+2(5x-5)=2 Think you can do that? Then, just plug in the value you get for x into y=5x-5 and solve for y.

OpenStudy (mckenzieandjesus):

so do i do 5+5 and -5+5?

OpenStudy (mckenzieandjesus):

5+5=10 -5+5=0

OpenStudy (anonymous):

a maybe

OpenStudy (haleyelizabeth2017):

no lol...

OpenStudy (haleyelizabeth2017):

my apologies, I'm getting help on my own math lol so... -4x+2(5x-5)=2 -4x+10x-10=2 Think you can take it from there?

OpenStudy (mckenzieandjesus):

i keep getting an answer thats not on my answer choices.. lol

OpenStudy (koikkara):

\(-5x+y=-5\) \(-4x+2y=2\) change (Multiply with - ) \(+4x-2y=-2\) Then, \(-5x+y=-5\) \(+4x-2y=-2\) solution: \(x-y=7\) option \(D\) \(2-(-5)=7\)

OpenStudy (koikkara):

@mckenzieandjesus Check it before you copy.

OpenStudy (koikkara):

\(NOTE:\) It's just a model... i have purposefully done wrong

OpenStudy (mckenzieandjesus):

huh?

OpenStudy (mckenzieandjesus):

so option d isnt right?

OpenStudy (koikkara):

Time up: \(-5x+y=-5\) \(+4x-2y=-2\) Solution: \(-x-y=-7\) substitute, Option D.(2,5) \(2 \) = \(x\) \(5 \) = \(y\) Substitute: \(-x-y=-7\) \(-2-5~=~7\) \(you~are~right~!!\) @mckenzieandjesus

OpenStudy (koikkara):

\(**error: ~ -2-5~=~-7\)

OpenStudy (anonymous):

U ALWAYS HAVE ERRORS \(BANG\)

OpenStudy (anonymous):

\(freddy laugh\)

OpenStudy (mckenzieandjesus):

?

OpenStudy (anonymous):

Five Nights At Freddy's never heard of it?